Solve the equation. (Enter your answers as a comma-separated list. Simplify your answers completely.)
6t^3 + 25t^2 = 66t

Answer by addingup(3677)

6t^3 + 25t^2 = 66t
6t^3 + 25t^2 -66 = 0
Use the quadratic formula, where a=6, b=25, and c=-66
t = [-b+-sqrt(b^2-4ac)]/2a
t = [-25+-sqrt(25^2-4(6)(-66)]/2(6)
t = [-25+-sqrt(625--15840]/12
Now we have two real roots. Let's simplify the radical
t = (-25+-47)/12
t = 22/12 or t = -72/12
t = 11/6 or t = -6
Or separated by a coma:
t = 11/6, t = -6
or if you want decimals, divide 11/6 and you get:
t = 1.8333, t = -6